OpenStudy (anonymous):

the perimeter of a retangle room is 166ft, find the length and width of the room if the length is 5 ft longer than twice the width

myininaya (myininaya):

2L+2W=166 L=5+2W

myininaya (myininaya):

2(5+2W)+2W=166 10+4W+2W=166 10+6W=166 6W=156 W=156/6

OpenStudy (anonymous):

let wifth = x then the length = 2x + 5 so 2(2x+5) + 2x = 166 solve for x and you have it

OpenStudy (anonymous):

are u ok with these answers ?